OpenStudy (anonymous):

A baseball is hit so that it travels straight upward after being struck by the bat. A fan observes that it takes 2.90 s for the ball to reach its maximum height. a) Find the ball's initial velocity. (b) Find the height it reaches

OpenStudy (jamesj):

If the ball has initial velocity u and is subject only to the force of gravity, what is the equation for the velocity v at time t? v = (something involving u, g and t) ... what is this expression?

OpenStudy (jamesj):

If you're really, really stuck, go back to first principles and definitions. Acceleration is by definition the change of velocity with respect to time. Here the velocity changes from u --> v over time t, and you know it is decelerating at a rate of -g. Therefore \[ Acceleration = \frac{change \ in \ velocity}{change \ in \ time} = \frac{v - u}{t} = -g \] Hence \[ v - u = -gt \] and \[ v = u - gt \] Now, that being the case, for what value of t is v = 0?